MARS DESERT RESEARCH STATION

Science Report – February 17th

Crew 175 Science Report 17Feb2017

Science Report SOL 5

Experiment: Seismometer
Person filling in the report: Mouadh Bouayad

We could deploy the seismometer this morning, during the EVA. It was quite hard to do, with the gloves, especially to put the seismometer correctly. It was nevertheless a success. We decided to put the Stanéo digitizer in a box in order to protect it, and we covered it with a piece of plastic. We have then tried to immobilize everything by putting stones here and there.
The weather on the rest of the sol was worrying: there was a lot of wind (impossible thing on Mars because of the low pressure on its surface). I fear that the structure might crumble. Unfortunately, we’ll have to wait until tomorrow to check this out, during the next EVA.

Experiment: Balloon
Person filling in the report: Simon Bouriat

We were first supposed to use the balloon during the EVA this morning. Unfortunately, the weather was not on our side. We still don’t have enough results to make any conclusions. We hope to use it again next week.
The crew engineer and I started working on our new project about emergency procedures. There are three main points in that study. First, we will study the emergency procedures observed during scuba diving. One of the big step in this study is the comparison between this sport’s dangers and EVA’s ones. Secondly, we will try to set up different emergency procedures depending on the issues and brief the rest of the crew about it. Finally, we will simulate unexpected health or material issues during EVAs and try to apply these procedures.

Simon Bouriat – HSO Crew 175

Science Report – February 16th

Science report

Experiment: Balloon
Person filling in the report: Simon Bouriat
The balloon went on the EVA yesterday. It worked really well even if the weather was a bit windy. Unfortunately, the platform broke when we stored it. Today, during the afternoon, I fixed it and I recoded the Arduino. The idea is to record pressure and temperature in the hab during the whole day tomorrow and to compare the results with the ones we got from the balloon flight (here after). It will give us the average characteristics on the ground.  We also want to calculate the height of the balloon thanks to our results. Tomorrow, during the EVA, we will try to tether the balloon with two ropes to avoid any rotation of the new platform.
I also tried to figure out how to build a radio relay. The main way in my opinion is to use one of the broken walkie-talkies.



Experiment: Seismometer
Person filling in the report: Mouâdh Bouayad
Today the battery was wrapped in an emergency blanket in order for it to be protected, especially from water. The assembly of the science equipment (ISAE-SUPAERO Seismometer and Staneo acquisition system) was assembled with the gloves in order to train for tomorrow’s EVA.

Experiment: Aquapad
Person filling in the report Arthur: Lillo
The Aquapad (CNES) experiment has started today, with water samples having been prepared for testing, it has been put in the science oven, the first results will be available tomorrow!

Experiment: EMUI
Person filling in the report: Louis Maller
During the EVA I wore the glasses to test different issues. Comfort: they are lighter and a lot less trying to wear than the previous version, this is a true improvement. The right branch of the glasses still hurt a bit, at the end of the EVA I was glad to take them off, maybe because of the cable attached to the external battery pulling down?
The glasses were attached to an external battery and I went through the entire EVA without any problems.
I was able to see the screen only if I managed to make some shade on the glasses, or if I was in a dark place, like in the shadows. The time widget was very visible.
Screencasting is still an issue for now as the available software either require internet to work, or require a cable…

Astronomy Report – February 16th

ASTRONOMY REPORT
SOL 3


NAME:  Mouadh Bouayad           CREW: 175
DATE: 02/15/2017

SKY CONDITIONS: clear.

WIND CONDITIONS: no wind

OBSERVATION START TIME: 10:30 pm

OBSERVATION END TIME: 11:00 pm

SUMMARY: I went to the observatory tonight, but the battery of the rotation motor was not charged enough. So it worked for few minutes and then stopped. I hope tonight it’ll be fully charged, and we’ll do observations (I plan first on observing Mars and Venus).

OBJECTS VIEWED: few stars naked eye again.

PROBLEMS ENCOUNTERED: The battery was still not charged enough.

Astronomy Report – February 15th

Crew 175 Astronomy Report 15Feb2017

ASTRONOMY REPORT
SOL 3
NAME:  Mouadh Bouayad   CREW: 175
DATE: 02/15/2017
SKY CONDITIONS: clear.
WIND CONDITIONS: no wind
OBSERVATION START TIME: 9:30 pm
OBSERVATION END TIME: 10:00 pm
SUMMARY: I replaced the battery, and only one of them lighted the LED. However, yesterday evening, there wasn’t enough charge to rotate the dome. We could hear a little noise while pushing the button, but nothing happened. We thus decided to leave it there, and to retry it today, and rotate the dome in order get more sunlight on the solar panel.
I returned to the observatory today at 4:15pm and TADAAAAH! It finally worked! I oriented the observatory north-west, for it to receive sunlight. I plan to make a long observation tonight! We could observe Mars and Venus if we go soon enough, and take some pictures of Orion nebulae. I hope we’ll be able to have great pictures!!

OBJECTS VIEWED: few stars naked eye.
PROBLEMS ENCOUNTERED: The battery was still not working, but it works now.

Science Report – February 15th

Crew 175 Science Report 15Feb2017

Science report Sol 3

Experiment : AR Glasses
Person filling in the report : Louis Maller

Today I continued working on the AR glasses, the issue with the screencast was solved, and I continued exploring the functionalities of the glasses, and inspiring myself from the needs around the Hab (engineering check for example), I am imagining more potential uses for the glasses. The realization might be an issue, but hopes are good. It’s interesting work, but I wish I had been able to do it earlier, which wasn’t possible since the glasses were in France and I was in Russia.

Science Report – February 14th

Science report SOL 2

Experiment : Seismometer
Person filling in the report : Mouadh Bouayad

I alone have been to the Science Dome today, at 4 :30 pm. It was real quick, as the commander, Arthur, asked me to be back at 5 :00pm. I brought a box with me, in order to put the instruments in it, and see if we can move it, already settled, during an EVA. I figured out that the battery is too heavy to put it in the box. I think we will move it independently from the other instruments, and we will plug the wires to it afterwards, in situ. The other instruments fit well in the box.

In order to protect the battery, particularly from the rain, I plan to use an emergency blanket to wrap the battery. I’ll leave a hole for the power wire to go through. We also would like to put a sign on the measurements site, in order to warn tourists not to touch the instruments. This would very unlikely happen, but let us be wary.

Experiment : Optinvent connected glasses
Person filling in the report : Louis Maller

I have been working with Mouadh on the Optinvent connected glasses. We are working on connecting the glasses to the computer screen using the dedicated application (Vysor), with mitigated success. We also installed the Tasker app and took some pictures with the integrated camera. We are also trying to find a technical solution in order to have visibility outdoors (looking into a strong light the screen is barely visible).
We are hoping to solve all these basic issues by tomorrow in order to be able to start the more complex and interesting tasks.

Experiment : Balloon
Person filling in the report : Simon Bouria

Concrete work on the solar balloon started today. With another member of the crew this morning, I built the platform of the balloon (made of cardboard boxes). The balloon can handle a two kilograms weight and has to support one or two GoPro cameras and the Arduino system. During the afternoon, I made the Java code of the Arduino to get the temperature, the pressure and the time. A few tests proved that the system is now functional. The solar balloon can now be taken for an EVA. We still have to take pictures and videos of the balloon, build a radio relay and prepare how and when we will use the balloon according to the weather. We still don’t know how high the balloon will go and how we are going to tether it. A complete protocol remains to be done to have a really efficient EVA. We also wondered if the second balloon will be built during an EVA or in the science dome.

Astronomy Report – February 14th

ASTRONOMY REPORT
SOL 1

NAME:  Mouadh Bouayad           CREW: 175
DATE: 02/14/2017

SKY CONDITIONS: pretty clear, even with few clouds here and there.

WIND CONDITIONS: no wind

OBSERVATION START TIME: 10:30 pm

OBSERVATION END TIME: 11:30 pm

SUMMARY: We managed to observe few objects in the sky. Unfortunately, I don’t know who’s the last one that used the observatory, but he or she forgot to turn of the switch back off, so the battery was drained of power. Therefore, we could open the observatory, but we couldn’t turn it right or left. We could however observe the moon, and Jupiter for few minutes. I hope that I will be able take some picture tomorrow.

OBJECTS VIEWED: Moon, Jupiter

PROBLEMS ENCOUNTERED: One of the batteries was dead so we couldn’t observe what we wanted.

Science Report – February 4th

Science Report:
Prepared by: Mamatha Maheshwarappa
Sol: 06
Earth Date: 02.04.2017
Title: Characterizing the transference of human commensal bacteria and developing zoning methodology for planetary protection
Project Advisor: Dr Lewis Dartnell, Professor of Science Communication, University of Westminster
Purpose: This research aims at using metagenomic analyses to assess the degree to which human-associated (commensal) bacteria could potentially contaminate Mars during a crewed mission to the surface. This will involve swabbing of interior surfaces within the MDRS habitat to characterize the commensal biota likely to be present in a crewed Mars mission, and collection of environmental soil samples from outside the MDRS airlock door and at increasing distances from the habitat (including a presumably uncontaminated site) in order to characterize transference of human commensal bacteria into the environment.
About the project: The internal samples (swab kits) are scheduled towards the end of the mission (but before anything has been cleaned or wiped down), so that the moist swabs with sampled bacteria spend as little time as possible before coming back to Earth for further analysis. As the external samples had no restrictions, it was scheduled on 4th Feb 2017.
With these soil samples, we are looking to see if human commensurate bacteria have escaped out of the MDRS habitat and into the surrounding area. Soil samples were taken (i.e. 3x Falcon tubes full of soil) immediately outside the airlock door, and then 1m, 2m, 5m and 10m from the airlock door. We are yet to collect at least a couple of samples from a location away from the MDRS that has not been visited before on previous missions as we need a ‘pristine’ sample of the desert soil that hopefully has not been contaminated with human commensurate bacteria escaping from the MDRS.
Equipment used:
⦁ 15 Falcon™ 50mL Conical Tubes (5 locations x 3 samples at each location)
⦁ Powder free Latex Examination Gloves
⦁ 70% Ethanol bottle
⦁ Personal Navigator

Protocol:
⦁ Squirt a small amount of 70% ethanol from the bottle and then rub it around the hands after putting on the gloves. Wait a few seconds for it to evaporate off.
⦁ Open one of the 50ml Falcon tubes and scoop up surface soil at the sampling location. Soil was collected at the surface across a wider area than digging a single hole really deep. Took total of three (3) 50ml Falcon tubes full of soil at each location, so that we get triplicate samples.
⦁ The indelible pen was used to label the outside of each Falcon tube with the sample number – e.g. 1a, 1b, 1c from the first site, 2a, 2b, 2c from the second location, etc.
⦁ The sample numbers and the location where they came from were noted and photos were taken for future reference. Also, the precise GPS coordinates were noted using personal navigator.
⦁ The collected soil samples were stored in the fridge.

Science Report – January 25th

Sol 10 Science Report – 3d printing project

25/01/17

3D Printing the groundbase for martian exploration
Crew Engineer

Needless to say that the first expedition to Mars will be difficult. It
will be even more difficult, not to say impossible, if nothing is prepared
in advance. The first crew won’t be able to book a hotel on Mars if that
hotel doesn’t even exist in the first place. They won’t be able to enjoy a
nice cup of tea and walk in T-shirt inside a habitat if that habitat hasn’t
been even built in the first place. My point here is that we need to
manufacture a few manned-infrastructures in advance and make sure these
maintain pressure conditions similar to those we can find on Earth while
providing reasonable thermal and radiation shielding.

Hopefully, with the 3d printing technologies currently available or in
development, the first habitats may be printed directly on Mars using
nothing else but INSITU resources and robots that would 3d-print and
assemble blocks together to design complex infrastructures.

The concept proposed and currently being studied at MDRS, is to manufacture
elementary blocks that can also contain water within their structure. The
water can be used not only for daily usage, but, can also provide extra
radiation shielding.

The first week at MDRS, we encountered several issues with the 3D printer
which didn’t allow us to print bricks but we managed to print 4 bricks over
the last days. Every brick takes 17h on average, and prints the outer shell
of the brick using PLA filament (plastic). For future studies, laser
sintering technology is suggested to simulate a real application on the
Martian soil. With the crew geologist Roy Naor, once the brick is printed,
we evaluate different types of soil that can be used within the brick to
strengthen it. Future tests are planned for the incoming days, and these
include, building a small infrastructure on an EVA, in order to prepare for
the next iteration of the concept.

The objective of this project is to give a first level of analysis so as to
lay down a first proof of concept.

Science Report – January 22nd

Sol 7 Science Report

GreenHab officer Rick Blake

One study being conducted by this crew rotation is designed by Israeli high school students. It involves investigating the differing ground colours on the hills around the MDRS and reporting our findings back to the students. Rather than a strict scientific aim, this experiment was mostly proposed to get students interested in STEM careers. The students don’t have any prior geological knowledge and it is hoped that by doing this experiment we will inspire their imagination and thirst to learn science. Our geologist, Roy, already knows the geological setting and history of the area to be tested, so any data generated will be purely for the students to interpret.
For this experiment, a transect up a hill was conducted during an EVA by our crew geologist, Roy, and myself. Samples were taken of the regolith and underlying rock at every change of colour on the ground. These samples were labelled and brought back to the Science Dome for further analysis. The samples were inspected for their general colour, reaction to acid, and, under a microscope; grain size, grain roundness, and any other interesting features. This data was recorded and will be relayed back to the students to interpret. Small sections of the samples will also be returned to the Davidson Institute for Science Education, Weizmann Institute, Israel for the students to further analyse.
For reference, it is known that the stratigraphy is part of the Brushy Basin member of the Morrison Formation, and it was formed in the late Jurassic in a fluvial lacustrine environment. The top of the stratigraphy is capped by Cretaceous geology, which ended up being the last sample collected on the transect.
The data recorded is as follows: